Systems Staff Meeting Notes
June 13, 2008
June 13, 2008
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Amy leaves tomorrow for Honduras.
- Logan notes that Missy is having a yard sale tomorrow.
- UF Exchange AD Meeting highlights - they are now in Phase Two & adopting new units. Materials Engineering adopted AD, but it was poorly planned (on their part). They advise advance testing and preparing users. Their IT folks did not have tier training. Issue with mailbox size - the size is based on current size. Attachments can blow it up. Will is going to contact Mike Conlon and Eric and try to work on a time frame for us to migrate to AD. Perhaps by the fall. Our current email addresses will still be useable. Departments now migrating include: Biomed, Law, CLAS.
- Today is the last day for VISA orders for this fiscal year.
- Loaner Laptop encryption - Will is testing. We may need to buy encryption software.
- Public workstation CJK - Logan to install the software today.
- MSL laptop return - Michael Howell has 32 boxes ready. Should go out on Monday.
- Office 2K7 - Amy made the link for download available on the Liaison Blog. It would be better to be able to push it to the desktop for Admin and Business Services, but there are difficulties with SMS. Logan to work on it next week.
- Systems training - Grover problems are being fixed. We will need to test and publicize it. Perhaps the cross-divisional meeting?
- The next Systems training may be Campus AD training.
- Reminder: the activities list is still on the table. Regular and future projects, day-to-day and long term as well.
Software Projects Status
- Wednesday we had a cluster server problem - Exchange and NT2R both became unavailable. We had to do a restart of the cluster services to get them back up. We're unsure of the cause. Logan had approved some updates that day. Maybe it was a switch problem? No real answer.
- Disaster Prep - haven't gotten too much further. We do have a one-man station set up in the server room. Discussed putting UPS on individual machines. Not many of the UPS's in the vault work. Logan took some - it took trying four to find one that was functional. May need to buy new ones and get rid of the old ones.
- SX280's - Dell techs on site June 23-27. Cynthia is coordinating a schedule for Liaisons to bring the machines to Systems so the tech can work on them here. They should be able to work on 16 per day.
- McAfee - Bobbie continues to work on it - verifying the push to the servers. Framework service needs to be installed.
- Greenstone backup - still holding.
- Service Pack 3 - issues? We have not pushed it yet. Amy has a thread about it on the Liaison Blog. Cynthia notes the laptops are waiting to install Service Pack 3 when updating.
- Campus AD - need to know the number of mailboxes by type. This will determine the cost.
- Shibboleth is going to happen. We will join.
- Grover training is done. Programmers are tweaking problems. We should be ready to test the roll out in a couple of weeks.
- DLC Tracking database - Tom has finished testing. Working on documentation for importer.
- Greenstone - one issue with the Development server - it was browse-able. It's been fixed.
- Greenstone - Mark made changes for UFDC. Bill approved the Fedora testing.
- Circulation Analysis - Winston received an email from Donna Alsbury at FCLA - she will pull a huge file and we will get going on this.
- Affiliation modification proposal - we're working around it right now, going with the 14-day accounts. Ultimately, however, a better method is desired.
- VIVO - Sara and Val are at a VIVO conference. They sent out a survey to name the project. Winston gave Logan a brief synopsis of the hardware requirements in order to get a monetary estimate of the cost to incorporate into the grant. Probably about $3500.
- Bryant Collection - Chris will start working on this as EAD moves forward.
Students and Department
- Greenstone Copybook diagnostics - they have been losing their network settings. Logan sent Randall some info on the problem.
Systems Security Management
- New work study student: Erasmo Garcia.
- Rob will continue through Summer B.
- Meeting held.